OpenStudy (anonymous):

e^x^2=e^12x times (1/(e^27)) what is the solution x=______________

OpenStudy (anonymous):

\[e ^{x ^{2}}=e ^{12x} \times \frac{ 1 }{ e ^{27} }\]

OpenStudy (anonymous):

right side = e ^(12x -27) do you agree with me?

OpenStudy (anonymous):

okay but what happens to the 1

OpenStudy (anonymous):

hey, it's multiplication, not addition.!! I multiply them.

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

good, now both sides is exponential of e , take e out, you have x^2 = 12x-27 ---> x^2 +12x+27=0. solve as usual to get x

OpenStudy (anonymous):

so the answer would be -3 and -9

OpenStudy (anonymous):

yes, sorry, you are right, negative

OpenStudy (anonymous):

oh nvm its x^2-12x+27 so you are actually right

OpenStudy (anonymous):

both us confused. do what you believe in.

OpenStudy (anonymous):

you move the 12x and -27 to the x^2 side so it makes 12x into -12x and -27 to 27 so making it x^2-12x+27 and the fraction would be (x-9)(x-3) and the zeros would be 3 and 9

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

lol got confuse when you put a pos. 12 instead of a neg 12. thank you!